Holly's Ham, Inc. sells hams during the major holiday seasons. During the current year 11,000 hams were sold resulting in $220,000 of sales revenue, $55,000 of variable costs, and $24,000 of fixed cost. The break-even point in sales dollars is: A. $320,000 B. $16,000 C. $32,000 D. $61,740

the cost of hot dog buns at National Coney Island would be classified as a: fixed cost, variable cost, labor cost, or none of the above?

An inaccurate cost function may result in:

A.

Evaluating a weak manager as providing strong performance

B.

Promoting a product that is actually less profitable than budgeted

C.

Predicting total costs that are too low

D.

All of the above
Answer #1

Break even sales = Fixed cost/Contribution margin ratio

Selling price per unit = 220,000/11,000 = 20

Variable cost per unit = 55,000/11,000 = 5

Contribution margin ratio = (20-5)/20 = 75%

= 24,000/75%

= 32,000

Option C

2. Variable costs
3. Option D